james is 5 years older than peter. if the product of their ages is 234, how old is peter?

Answer by ikleyn(52790)

.

There are two ways to solve this problem.


    1)  You may find decomposition of the number 234 into the product of two positive integer numbers 

        with the difference 5 between them  (MENTAL solution (!) )


            234 = 117*2 = 13*9*2 = 13*(9*2) = 13*18,  which is exactly what you need,


OR


    2)  you may solve the quadratic equation


            x*(x+5) = 234.


    Surely, in this way, you will get the same answer.

ANSWER.   The younger,  Peter,  is  13  years old.